# GOLLuM: Gaussian Process Optimized LLMs – Reframing LLM Finetuning through Bayesian Optimization

### v2

@@ Abstract
Large Language Models (LLMs) can encode complex relationships in their latent spaces, yet harnessing them for optimization under uncertainty remains challenging. We address this gap with a novel architecture that reframes LLM finetuning as Gaussian process (GP) marginal likelihood optimization via deep kernel methods. We introduce LLM-based deep kernels, jointly optimized with GPs to preserve the benefits of both – LLMs to provide a rich and flexible input space for Bayesian optimization and – GPs to model this space with predictive uncertainty for more efficient sampling. Applied to Buchwald-Hartwig reaction optimization, our method nearly doubles the discovery rate of high-performing reactions compared to static LLM embeddings (from 24% to 43% coverage of the top 5% reactions in just 50 optimization iterations). We also observe a 14% improvement over domain-specific representations without requiring specialized features. Extensive empirical evaluation across 19 benchmarks – ranging from general chemistry to reaction and molecular property optimization – demonstrates our method’s robustness, generality, and consistent improvements across: (1) tasks, (2) LLM architectures (encoder, decoder, encoder-decoder), (3) pretraining domains (chemistry-related or general-purpose) and (4) hyperparameter settings (tuned once on a single dataset). Finally, we explain these improvements: joint LLM-GP optimization through marginal likelihood implicitly performs contrastive learning, aligning representations to produce (1) better-structured embedding spaces, (2) improved uncertainty calibration, and (3) more efficient sampling – without requiring any external loss. This work provides both practical advances in sample-efficient optimization and insights into what makes effective Bayesian optimization.
1 Introduction
Large Language Models (LLMs) have demonstrated remarkable capabilities in natural language understanding and generation 1, 2, 3, 4. Their success stems from an ability to learn rich representations of text that capture subtle patterns, relationships, and domain-specific knowledge 5, 6. This representational power has naturally led to growing interest in adapting LLMs beyond general language tasks to specialized domains – from scientific discovery to reasoning tasks 7, 8, 9, 10, 11. However, despite their expressive capabilities, LLMs exhibit fundamental limitations in reliability. Even in their primary domain of text generation, they can produce overconfident yet factually incorrect outputs through hallucination 12, 13, 14. In high-stakes fields like drug discovery, materials design, or automated reasoning, such failures carry critical real-world risks, making principled uncertainty quantification essential 15, 16, 17, 18, 19, 20, 21.
In this context, Bayesian optimization (BO) 22, 23, 24 has emerged as a powerful strategy for optimizing expensive-to-evaluate functions by efficiently balancing exploration and exploitation 25, 26, 27, 28, 29. BO typically employs Gaussian Processes 30 (GPs) due to their principled uncertainty estimates, interpretable confidence bounds and well-calibrated predictions even in low-data regimes. This property of GPs is particularly valuable in domains that demand sample-efficient optimization of complex (e.g., analytically intractable) objectives – as is often the case in chemistry. Recent works have begun exploring LLMs for Bayesian optimization, either as feature extractors for surrogate models 31 or through sequential LLM finetuning with post-hoc uncertainty quantification 32. While promising, these methods either underutilize LLMs’ adaptation capabilities or decouple predictive performance from uncertainty estimation, limiting their optimization effectiveness.
We introduce GOLLuM (Gaussian Process Optimized LLMs), a framework that seamlessly integrates LLMs into the GP architecture through Deep Kernel Learning (DKL) 33, 34. Rather than using LLMs as sophisticated encoding tools (via embeddings) or lookup-tables (via prompting), our approach directly employs the GP marginal likelihood as LLM finetuning objective. The resulting method provides a bidirectional feedback loop: the GP guides updates to LLM weights to produce more effective embeddings, which in turn enhance the GP’s probabilistic modeling. Through GP marginal likelihood optimization, the LLM embeddings adapt to follow a key principle: points with similar function values should be close in the embedding space, aligning with the similarity structure imposed by the GP kernel. This strategy inherently induces a contrastive learning effect 35 enabling the model to naturally organize the design space points into distinct regions – the good (high objective function values), the bad (low objectives), and the ugly  [^1] Reference to the 1966 film The Good, the Bad and the Ugly, directed by Sergio Leone.. The clear separation in the latent space improves the optimization performance by enabling more effective exploration while the joint training helps maintain well-calibrated uncertainty estimates, despite potential kernel misspecifications.
Our framework is agnostic to the specific LLM architecture or its pretraining, effectively adapting any general-purpose LLM into a powerful optimization tool with rigorous uncertainty quantification. We evaluate our approach within the chemistry domain, where expensive evaluations and vast design spaces pose crucial bottlenecks to progress in drug discovery and materials science 29, 36. In this context, our method’s ability to efficiently navigate complex optimization spaces with principled uncertainty guidance has the potential to accelerate scientific discovery.
Our key contributions include:
  1. 1.
    GOLLuM: The first jointly trained LLM-based deep kernel GP architecture. We introduce the first end-to-end framework that integrates LLM capabilities with Gaussian processes via marginal likelihood, enabling principled Bayesian optimization. We demonstrate and formalize how GP marginal likelihood optimization induces contrastive structure in the embedding space – separating regions by performance – without any explicit contrastive loss. This unified approach provides an alternative to domain-specialized models or representations while producing interpretable latent space structures.
  2. 2.
    Empirical insights into representation factors that enable successful high-dimensional BO. Through systematic analysis of 14 fixed LLM and chemistry-specific representations, we quantify how representation structure influences BO success. We find that optimization success strongly correlates ($r=0.92$) with a normalized smoothness metric capturing the alignment between the GP’s inductive bias and the structure of the representation space. Representations that support smooth yet calibrated surrogate fits enable more principled exploration and lead to better BO outcomes.
  3. 3.
    Robust generalization across diverse chemical tasks. Our approach demonstrates consistently strong performance across 19 diverse chemistry benchmarks, using hyperparameters tuned on a single dataset. Compared to chemistry-specific representations, fixed LLM embeddings, and disjoint finetuning approaches, our method achieves superior exploration of chemical spaces, better sample efficiency, and generalization. This conclusion is supported by more than 8,000 experiments.

### v3

# Large language models as uncertainty-calibrated optimizers for experimental discovery

@@ Abstract
Scientific discovery increasingly depends on efficient experimental optimization to navigate vast design spaces under time and resource constraints. Traditional approaches often require extensive domain expertise and feature engineering. While large language models, with their vast scientific knowledge, circumvent the feature engineering limitations, they lack the calibrated uncertainty estimates required for high-stakes decision making. Hence, current optimization methods force a choice between domain knowledge and reliability, with no principled approach that affords both. In this work, we show that training language models through the uncertainty-aware objectives of traditional optimization methods enables their use as reliable optimizers guided by natural language. By teaching LLMs from experimental outcomes under uncertainty, we transform their overconfidence from a fundamental limitation into a precise calibration mechanism. Applied to Buchwald-Hartwig reactions, a cornerstone of pharmaceutical synthesis, our method nearly doubles the discovery rate of high-yielding reaction conditions – 24% to 43% in 50 experimental iterations starting from 10 unsuccessful conditions. Across 19 diverse optimization problems spanning organic synthesis, materials science and catalysis, process chemistry, and molecular design, our approach ranks first on average, establishing a new paradigm for reliable, uncertainty-guided optimization with LLMs. Our approach can accelerate discovery by lowering the barrier to using powerful optimization methods, replacing the need for domain-specific feature engineering with more accessible natural language interfaces. These findings highlight that ensuring reliability through principled uncertainty quantification is critical for realizing the full potential of AI-guided experimentation.
1 Main
Artificial intelligence can now generate millions of novel drug candidates, material structures, or chemical reactions in a matter of hours, expanding an already vast combinatorial landscape that has long challenged experimental science 1, 2, 3. However, the fundamental bottleneck remains: deciding which candidates to evaluate next. This explosion of possibilities demands systematic, sample-efficient approaches to navigate vast design spaces under time and resource constraints 4, 5, 6, 7.
Established experimental optimization methods, such as Bayesian optimization (BO) 8, 9, 10, 11, 12, 13, 14, address this problem through a probabilistic lens. BO relies on surrogate models, typically Gaussian processes (GPs) 15, to provide experimental predictions together with uncertainties, balancing exploration and exploitation 16, 17, 18. In practice, however, BO typically depends on domain-specific feature representations 19, 20, 21, 22. Optimizing a chemical reaction, for example, requires first translating molecular structures, catalytic mechanisms, and reaction pathways into descriptors that encode chemical intuition 23, 24. When shifting domains from organic synthesis to materials discovery, the time-consuming feature engineering must be repeated, as descriptors for one field do not apply to the other. Reaction fingerprints cannot characterize crystal structures, catalyst optimization bears little resemblance to drug design, and even within catalysis, descriptors optimized for a metal-ligand system may not be effectively transferred to others 25, 26, 27. Although BO provides a general approach, each new domain effectively starts from scratch with little performance guarantee, creating barriers to adoption and limiting the transferability of expertise 28.
In contrast, large language models (LLMs) have emerged as generalist models that transcend domain boundaries through rich prior knowledge acquired from their vast pretraining data 29, 30, 31, 32. The potential use of LLMs for experimental optimization has attracted considerable interest, yet adoption remains constrained by a fundamental flaw: unreliability. LLMs hallucinate, producing outputs that sound confident but are factually incorrect 33, 34, 35. Although this limitation frustrates users in everyday applications, it has severe consequences in experimental science. Hallucinated experimental suggestions waste resources, delay research timelines, and compromise safety protocols in automated laboratory settings 36, 37. Unsurprisingly, their integration into high-stakes experimental workflows remains impractical without rigorous control 38, 39.
The merging of the two paradigms, BO and LLMs, has therefore gained significant traction 40, 41, 42, 43, 44, 45. Yet existing approaches suffer because they treat language models and probabilistic optimization separately. We argue instead that the resolution lies in recognizing they should exist as a unified system. Here, we introduce GOLLuM (Gaussian Process Optimized LLMs) as the first framework to train language models through the same probabilistic objective that makes Bayesian optimization reliable. Rather than positioning LLMs as direct experimental optimizers, where they struggle with hallucination and heuristic search, we demonstrate how their probabilistic integration within established optimization mechanisms enables principled decision-making from natural language.
The framework establishes a joint optimization mechanism between LLM embeddings and probabilistic modeling. Given textual descriptions of prior experiments – reaction conditions, synthesis procedures, process parameters – the LLM transforms these into embeddings serving as inputs to a GP surrogate. Critically, the learning signals flow back from the GP to update the LLM. Through this bidirectional mechanism, the LLM learns to encode experimental performance rather than textual similarity, organizing its latent space into distinct regions of high- and low-performing experiments.
This learned organization reveals structure in the experimental landscape while providing interpretability. In Buchwald-Hartwig cross-coupling reactions, fundamental to medicinal chemistry, the transformed representations expose genuine chemical patterns explaining reaction performance. The model automatically groups reactions by their key chemical components, clustering iodide-based reagents in high-yield regions while separating them from low-performing bromides and chlorides. This separation mirrors known reactivity trends, offering rationale for why recommended conditions should work – yet it emerges entirely from the experimental data, without explicit chemical instruction. By exploiting these discovered patterns, GOLLuM nearly doubles the discovery rate of high-performing conditions, reaching 43% of the highest ranking conditions within 50 iterations compared to 24% with static embeddings. This improvement extends to other reaction classes, such as Suzuki-Miyaura cross couplings and decarboxylative arylations, while still outperforming domain-specialized descriptors. Most remarkably, the approach transfers across domains: from materials synthesis and process engineering to molecular design. Across 19 diverse benchmarks covering different optimization paradigms (combinatorial, mixed variables, continuous), GOLLuM ranks first on average using fixed hyperparameters (tuned once on a single dataset) starting from 10 failed experiments (below the median).
Our results suggest that uncertainty, rather than being LLMs’ fatal flaw for experimental science, can drive their transformation into reliable optimizers. By training language models together with the GP surrogates, we enable what neither paradigm could achieve alone: the generality to handle any experimental procedures described in natural language with the reliability required for real experimental campaigns. This principle of uncertainty-driven adaptation suggests a broader paradigm for deploying AI in high-stakes domains where calibrated confidence matters as much as capability.
Figure 1: Language models as uncertainty-calibrated optimizers a, Experimental optimization challenges. Scientists face the universal problem of efficiently exploring vast design spaces under time and cost constraints. b, Current approaches impose a trade-off between accessibility and reliability. Bayesian optimization provides principled uncertainty calibration and sample efficiency but requires domain expertise. LLMs offer natural language interfaces and prior knowledge but lack reliability. c, GOLLuM optimization loop. Previous experimental results, described in natural language, feed into joint optimization of both GP and LLM components. Likelihood gradients from the GP finetune the LLM, progressively smoothening the optimization landscape and organizing the search space into distinct regions: "the good", "the bad" (and the ugly) of high and low performance. This structured representation enables easier selection for the next experiment, which becomes a new training point that continues the loop. d, Key results. GOLLuM ranks first averaged across 19 cross-domain tasks spanning materials, reactions, molecules, and processes, enabling sample-efficient discovery from natural language inputs while surpassing both specialized descriptors, domain-pretrained models, and proprietary LLMs.
2 Results
2.1 Large language models are unreliable as direct optimizers
Using large language models as direct, prompt-based optimizers is an appealing strategy for leveraging their vast prior knowledge in scientific discovery 46, 41, 47, 48. In such scenarios, the LLM directly proposes experimental conditions based on the prompt containing details of the design space, previous experiments, and the optimization goal. However, this application relies on the model’s black-box patterns and ignores known limitations, including overconfidence and a lack of principled uncertainty calibration.
We first evaluated this approach for experimental optimization by prompting LLMs to act as expert chemical optimizers (Figure 2a). We analyzed results of several state-of-the-art proprietary, closed-source LLMs to suggest new conditions for Buchwald-Hartwig 49 reactions, providing them with the parameter space and a history of previous experiments. This approach, however, proved fundamentally unreliable for scientific discovery.
A quantitative analysis of the models’ responses revealed a high failure rate across all tested LLMs (10% to $\sim$80% in Figure 2c, Response type). A significant fraction of suggestions were invalid due to hallucinated chemical structures or proposals outside the defined parameter space. Furthermore, models frequently suggested duplicate conditions, failed to generate parsable outputs or tried to prematurely end the optimization campaign. These failures stem from the inherent limitations of using LLMs as direct optimizers: their lack of principled uncertainty calibration and their reliance on heuristic search rather than a principled probabilistic strategy.
While LLMs offer integrated reasoning and contextual analysis of the optimization problem, there are currently no clear mechanisms to evaluate or mitigate their overconfidence or outright hallucinations in a quantifiable manner. This black-box reasoning process makes them unsuitable for high-stakes experimental decision-making and underscores the need for a framework that integrates their domain knowledge within a reliable, uncertainty-aware optimization loop. Critically, when benchmarked against traditional Bayesian optimization, the LLM-driven approaches substantially underperformed, failing to consistently identify the top-performing reactions, and trailing behind our proposed GOLLuM method (Figure 2c, Top 5% Coverage).
The failure of direct prompting motivates a more principled approach. Besides their generative abilities, LLMs offer their scientific knowledge in a more constrained manner: as sophisticated feature extractors for a traditional Bayesian optimization framework. We next investigate this strategy by using static LLM embeddings as chemical descriptors.
Figure 2: From LLM prompting to LLMs as principled scientific optimizers. a, Querying LLMs for optimization. Given the parameter space and previous experiments, the LLM proposes next suggestions. This method relies on heuristic search and suffers from a lack of uncertainty calibration, which leads to invalid suggestions, hallucinations, and frequent failures. b, GOLLuM framework integrates an LLM and a GP in a joint optimization loop. Gradients from the GP’s probabilistic objective finetune the LLM, creating an interpretable search space that enables principled, uncertainty-aware, and sample-efficient optimization. c, Quantitative results on the Buchwald-Hartwig reactions benchmark. (Left) The combinatorial design space of 3955 reactions across 5 distinct chemical products, corresponding to $\sim$800 unique conditions per product. (Middle) The analysis of response types from several proprietary LLMs reveals failure rates, with a high fraction of invalid or duplicated suggestions. We executed 5 independent optimization seeds (API-cost limited) per product. Each run cold-started with 10 initial low-performing reactions and repeatedly queried the model until reaching a quota of 50 unique, in-space suggestions. The stacked bars aggregate the raw outputs from all queries issued while reaching that quota across 25 runs (5 products x 5 seeds) per model (>>1250 queries). Categories: Valid, Duplicate (already proposed in the same run), Invalid (outside the parameter space), and LLM failure (format/parse/timeout). (Right) A comparison of optimization performance ("Top 5% Coverage") shows that direct LLM approaches substantially underperform traditional Bayesian optimization (with reaction fingerprints 50) and our proposed GOLLuM method using GP-finetuned general-purpose open-source LLMs (T5 51). Bars show mean and standard error. Performance comparison excludes Claude-3.5 Haiku (did not reach the quota of 50 valid in-space suggestions).
2.2 Static LLM embeddings underperform as chemical descriptors
A critical requirement for scientific optimization is a representation that captures the relationship between experimental parameters and outcomes 22. Traditional chemical descriptors rely on structured data, leaving the vast majority of scientific knowledge constrained in unstructured language such as patents, publications, and even chemical intuition. Here, we evaluate whether LLMs can translate the chemical knowledge from text into data-driven descriptors for Bayesian optimization tasks.
On the same benchmark set of Buchwald–Hartwig reactions 49, we observed that static LLM representations consistently underperformed specialized chemical descriptors (Figure 3a). In 50 optimization trials initiated from a realistic cold-start scenario with 10 low-performing experiments, LLM embeddings helped discover only 15–26% of the highest-yielding reactions (5th percentile threshold), trailing behind the 38% achieved by specialized chemical fingerprints (DRFP 50). OpenAI’s embeddings 52, while reaching the highest score among the LLM embeddings, still failed to close this performance gap.
The simple lack of chemical knowledge does not explain these differences. Chemistry-specialized language models, such as T5Chem 53, underperformed general-purpose variants, achieving a 23% discovery rate despite extensive pretraining on chemical tasks. The model’s dependence on input format revealed underlying patterns that explain the disparity in results. Changing T5Chem’s input from an experimental procedure to the Simplified Molecular Input Line Entry System (SMILES) format 54, 55, 56 improved its discovery rate by 44%. This improvement is not surprising given the model’s inherent familiarity with the SMILES notation from its pretraining. Such format dependency indicates the core issue is not the knowledge the model contains, but rather its ability to translate that knowledge into the embedding space.
These results highlight the limitations of static LLM embeddings for experimental guidance. Without direct adaptation to the optimization task, they force a trade-off between the convenience of natural language and the optimization performance of structured descriptors.
Figure 3: Representation geometry enables efficient Bayesian optimization a, BO performance with fixed LLM features as input to GP. We benchmark three encoder models (ModernBERT 57, UAE 58, MXBAI 59), three encoder–decoders (Instructor 60, T5 51 and its chemistry-related variant T5Chem 53) and four decoder families (Llama 61, 62, Mistral 62, Qwen 63, 64, and OpenAI 52). We show average discovery of high-impact regions of the design space as the percentage of the top 5% reactions found during the optimization, aggregated across all five Buchwald-Hartwig reactions. Chemistry-related representations include T5Chem-SMILES53, a pretrained chemistry-related LLM with SMILES input, and DFRP50, a reaction fingerprint. b, Comparative analysis of GP-based LLM finetuning. The finetuned models are arranged by the overall performance and relative improvements to their base (fixed embeddings) LLM-GP variants. Chemistry-related baselines (previous best) included for comparison. All results show mean and standard error across 20 independent seeds, with each optimization run starting from 10 initial points (random below median) and selecting subsequent points via acquisition function maximization over the enumerated design space. c, Data representations and their success rates in BO. BO performance correlates with GP smoothness, measured as the ratio of learned lengthscale to average pairwise embedding distance. Higher ratios indicate smoother GP fits, implying the representation supports meaningful generalization across the design space.
2.3 Representation geometry determines optimization success
The representation geometry determines optimization success better than predictive accuracy of a surrogate model (Figure 3c). Although regression metrics such as $R^{2}$ explain how well the surrogate fits the objective function, they do not offer a clear answer as to what dictates the optimization performance. Instead, we identify a simple metric based on representation geometry that correlates better with optimization success.
The ratio of the GP’s learned lengthscale ($\ell$) to the average pairwise distance between points in the search space ($\overline{d}$) robustly predicts optimization success. The lengthscale defines the distance in the input space over which the model expects output values to be correlated. A higher $\ell/\overline{d}$ indicates a smoother, more GP-friendly embedding space where the model can generalize over larger area because the embeddings position similar outcome experiments together. Across all fourteen representations, this geometric measure correlates strongly with BO performance ($r=0.92$, versus $R^{2}:r=0.78$, weighted $R^{2}:r=0.82$, see Figure 10 in SI). By contrast, representations with a low smoothness ratio create rough, difficult-to-model objective landscapes with potentially many local minima.
These findings underscore a well-structured representation as a more important factor for successful optimization than the raw predictive accuracy itself. If a GP surrogate with a stationary kernel (where the relationship between two points depends only on their distance) assumes a smooth objective function, the underlying representation should support it. Specialized chemical descriptors (DRFP 50) instill some organization in the latent space, resulting in a higher ratio ($\ell/\overline{d}\approx 1.8$) and better performance, while the static LLM embeddings exhibit low ratios, resulting in rough, difficult-to-model objective landscapes. The failure of static LLM embeddings is therefore a failure of alignment between the representation’s geometry and the surrogate model’s requirements, motivating a method that can actively learn to resolve this incompatibility.
2.4 Uncertainty-guided LLM finetuning enables efficient optimization
To resolve the misalignment between the representation and the surrogate model, the LLM cannot be a static feature extractor; it must learn to adapt. We therefore formalize pretrained LLMs as learnable feature extractors, "taught" by the GP’s own probabilistic objective. This approach transforms the LLM from a static component into a dynamic one, tailoring its vast pretrained knowledge to the optimization task at hand.
In the same Buchwald–Hartwig reactions benchmark as above, our joint GP–LLM training substantially improved optimization efficiency (Figure 3b). An encoder-decoder Google T5 51 model increased the discovery rate of high-yielding conditions from 24% to 43%, with a 79% relative improvement compared to its static representation. This increase surpassed specialized fingerprints, achieving the new state-of-the-art.
Similar improvements held across other architecture types. The same approach elevated decoder-only models (61% relative improvement for Qwen-7B 63), encoder types (42% for ModernBERT 57), and raised the score of closed source OpenAI 52 embeddings (by 31%) without the access to the base model’s weights. These systematic improvements across major LLM architecture types suggest a general principle rather than an isolated finding: that task-specific adaptation is more important than domain-specific pretraining. Indeed, general-purpose LLMs with no chemistry-specific pretraining (T5 51, Qwen-7B 63) surpassed both chemistry specialized features (DRFP 50) and specialized chemistry models (T5Chem 53).
The mechanism behind the sample-efficiency lies in the training objective. By finetuning the LLM with the GP’s marginal likelihood, we merge both components under a shared probabilistic objective that inherently balances data fit with uncertainty estimates. This training paradigm repurposes the uncertainty into an LLM learning signal that guides the reorganization of its latent space. Traditional LLM finetuning has often required thousands of training examples. Here we instead finetune large pretrained language models with 10 initial points (adding one example per iteration in 50 sequential trials) adapting their latent spaces into smooth and calibrated representations for the optimization task.
These results indicate that uncertainty-driven finetuning can unlock task-specific effectiveness from generic LLMs, bypassing the need for specialized model training in each new domain. By unifying natural language with probabilistic calibration, our approach enables the accessibility of language models and the reliability of Bayesian optimization within a single framework.
2.5 Emergent chemical organization supports interpretable optimization
[…]
Beyond improved performance, the finetuned LLM’s latent space exhibits emergent organization of chemical knowledge that offers interpretability. A persistent challenge limiting the adoption of AI in experimental science is the ’black box’ problem: even if the model’s suggestions are effective, they often lack a clear scientific rationale, forcing researchers to follow them blindly rather than with understanding. This prevents the generation of new scientific insights and undermines trust in the optimization process.
Our framework transforms this paradigm by making the model’s reasoning transparent, as uncertainty-guided finetuning fundamentally reorganizes the LLM’s latent space into an interpretable map of chemical reality (Figure 4c). GOLLuM autonomously rediscovers chemical patterns directly from experimental data, organizing the search space into meaningful regions of high and low performance. The reasoning for its suggestions, therefore, lies within the geometry of the latent space itself.
During optimization on the Buchwald-Hartwig tasks, the model automatically learns to cluster experiments according to key chemical features and their outcomes (Figure 4c). For instance, experiments using high-reactivity iodide-based aryl halides cluster in a region of the latent space associated with high yields, while those using less reactive chlorides occupy a different, lower-performance region. This data-driven grouping mirrors well-known reactivity trends in cross-coupling chemistry, despite the model never being explicitly taught these relationships. This organization emerges progressively, with structured patterns becoming observable by the 10th optimization iteration as the GP marginal likelihood objective acts as an implicit contrastive loss, separating successful and unsuccessful experiments (Figure 4d).
This emergent structure is functionally critical to the optimizer’s efficiency. By grouping successful experiments and distancing failures, the model creates a smoother, more coherent landscape for the GP, enabling more informed and sample-efficient exploration. This clustering transforms the model from a black box into a collaborative tool for scientific discovery. The organized latent space provides a direct rationale for the model’s suggestions, facilitating understanding of why a proposal is likely to succeed based on its proximity to known positive outcomes. The location of a proposed experiment in this latent space carries potentially beneficial information even before running the experiment, enabling new hypotheses and interaction with the optimization process. This interpretability turns a passive cycle of suggestions into an active dialogue between human intuition and machine intelligence, allowing scientists to gain new insights during the optimization rather than solely following outputs.
2.6 Broad generalizability across diverse chemical domains
[…]
The central premise of language model-based optimization is the potential for a single framework to leverage prior knowledge across diverse scientific problems without domain-specific feature engineering. Traditional optimization approaches rely heavily on expert-built descriptors tailored to each domain. We hypothesize whether the expressive power of language models under probabilistic alignment could capture the underlying physics and chemistry encoded in textual descriptions, eliminating the need for specialized representations by navigating the optimization from text.
To evaluate this capability, we benchmarked GOLLuM across 19 optimization tasks spanning organic synthesis, materials science, process engineering, and molecular design. These tasks encompassed diverse problem structures, including combinatorial search over discrete molecular spaces, continuous optimization of process conditions, and mixed-variable problems combining both discrete and continuous parameters. Using a single set of fixed hyperparameters without any task-specific modifications, we benchmarked GOLLuM against relevant baselines including specialized feature-based methods, fixed LLM embeddings, and supervised uncertainty approaches (summarized results in Figure 5c).
GOLLuM ranked first on average across all 19 benchmarks, consistently identifying optimal or near-optimal solutions with high sample efficiency. The framework maintained robust performance despite cold start initialization and achieved substantial quantitative improvements over competing paradigms ( 8% over specialized features, 23% over fixed LLM embeddings, and 114% over supervised uncertainty methods, Figure 5a).
The improvement over specialized features represents a particularly relevant achievement, as these representations typically encode decades of domain expertise and have historically enabled effective optimization in their respective fields. For example, molecular fingerprints capture specific structural motifs known to correlate with chemical properties, while reaction descriptors encode core reactivity patterns. That a language model-based approach can exceed these carefully crafted features suggests that simple textual descriptions of experiments contain richer, more generalizable information than previously recognized.
The 23% improvement over fixed LLM embeddings validates our hypothesis that optimization-aware finetuning benefits over static representations. Moreover, this approach reduced format dependency that limited fixed embeddings by enabling both the general T5 51 and T5Chem 53 to achieve comparable performance across both SMILES strings and natural language procedure descriptions (Figure 5b). This format robustness facilitates applicability to optimization problems with multiple data representations. In certain cases, however, particularly those involving small numerical values (such as HPLC flow rates of 0.02 mL), finetuning appeared to degrade performance relative to fixed embeddings. This limitation may arise when numerical precision requirements exceed the differentiable information content available in natural language descriptions, highlighting an important boundary condition for text-based optimization approaches.
The highest improvement (114%) came from integrating uncertainty quantification directly into the training process through Gaussian process-based finetuning, rather than treating uncertainty estimation as a post-hoc learning task. This design ensures that both predictive capacity and uncertainty calibration improve simultaneously, whereas approaches that prioritize the former over the latter may compromise reliability, effectively training models to be good guessers rather than calibrated reasoners 72. (Table 1).
Model $R^{2}$ $\uparrow$ NLPD $\downarrow$
[…]
GP+DRFP $0.4\pm 0.3$ $4.9\pm 5.0$
Bochem.+T5 $0.3\pm 0.2$ $38.3\pm 63.3$
Bochem.+T5Chem $0.5\pm 0.1$ $49.2\pm 62.3$
LAPEFT+T5 $0.4\pm 0.1$ $103.2\pm 203.4$
LAPEFT+T5Chem $0.5\pm 0.1$ $357.9\pm 1016.6$
[…]
3 Discussion
In this work, we propose how to resolve the critical trade-off between the domain knowledge of large language models and the principled reliability of Bayesian optimization. By repurposing the uncertainty from a Gaussian process as a direct training signal, we transform a generalist language model into a sample-efficient, domain-aware scientific optimizer. This transformation enables what neither paradigm could achieve alone: the generality of a text-based interface with the reliability required for experimental campaigns.
Our empirical evaluations demonstrate how finetuning language models with Gaussian process uncertainty, consistently matches and outperforms specialized, expert-engineered features in guiding scientific optimization. This robust mechanism allows the framework to generalize effectively across 19 disparate scientific domains without task-specific modifications. Such cross-domain success stems from an emergent understanding of the underlying optimization landscape, a principle quantified by the strong correlation between our geometric smoothness metric ($\ell/\overline{d}$) and search efficiency. This combination of a unified, text-based approach with a reliable probabilistic formulation makes it a powerful tool for accelerating discovery in self-driving laboratories and human-AI collaborative experiments.
While our approach demonstrates broad applicability across the chemical sciences, we acknowledge several limitations that suggest avenues for future research. The cubic scaling of standard GPs may pose a challenge for optimization campaigns with thousands of experiments, suggesting a future need to integrate scalable GP approximations 73, 74. Furthermore, while text provides a powerful universal interface, its effectiveness may be reduced in domains dominated by complex, high-dimensional data lacking a clear linguistic structure, such as protein conformations or crystal structure. Future work could explore hybrid models that combine our framework with specialized encoders for such data types.
This work suggests that a path to more capable and reliable LLMs for science lies in grounding the models in principled probabilistic reasoning. Crucially, our findings reframe uncertainty from a flaw to be mitigated, to a rich signal making language models more reliable and aligned with scientific goals. The ability to use natural language to optimize a chemical reaction, a material synthesis protocol, or a molecular design lowers the barrier to entry for researchers across all scientific disciplines. These results suggest a new standard for AI-guided experimental design: an accessible, interpretable, and reliable toolkit applicable within experimental science. By grounding LLMs in the principled probabilistic objectives, we fundamentally shift their learning goals from making best predictions to providing well-calibrated ones. This approach suggests a different paradigm for achieving specialization, not through more data, but through richer, uncertainty-guided information.